Fermi National Laboratory


Accelerator Update

Friday January 23
The day shift began with the Tevatron (TeV) recovering from a quench and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 10:30 AM.

At 1:51 PM, Operations established store 3187 with an initial luminosity of 12.64E30. Experts made eleven transfers in all, but two were lost. This was the first successful store established with antiprotons coming from the Recycler. The Antiproton Source (Pbar) resumed stacking.

A Main Injector (MI) expert completed a slipstacking study at 5:15 PM.

About 8:40 PM, TeV experts began an end-of-store study.

Operations terminated store 3187 at 9:38 PM.

At 10:15 PM, Operations put the TeV into shot setup.

Saturday January 24
The midnight shift began with the TeV in shot setup and with MiniBooNE taking beam.

At 12:35 AM, Operations established store 3189 with an initial luminosity of 45.51E30. Pbar resumed stacking.

A Linac RF station (KRF5) tripped off at 2:02 AM due to phase errors. Stacking resumed at 2:25 AM.

A Switchyard expert began a study period at 9:46 AM.

At 12:08 PM, a possible septum power supply problem developed with the Pbar Debuncher. Operators called in an expert. The expert said the power supply was okay but there was an open indication in the tunnel that would require an access.

A NuMI expert began a study period at 5:43 PM.

Sunday January 25
The midnight shift began with Operations monitoring store 3189, with NuMI studies, with MiniBooNE taking beam, and with stacking on hold until its septum problem (D:ISEPV) could be investigated.

At 1:48 AM, a Booster power supply (ORBMP) tripped off on an overcurrent indication. Operators called in an expert who discovered a bad capacitor. He had the power supply repaired and operating normally by 8:31 AM.

Resumed sending beam to MiniBooNE at 8:34 AM.

Established beam to the MI dump at 8:44 AM.

The Booster ORBMP power supply tripped again on overcurrent at 9:09 AM. The expert asked for more help and together they found and repaired a shorted SCR bridge by 11:49 AM.

At 11:53 AM, Operations established beam to MiniBooNE.

Operations terminated store 3189 at 12:06 PM and put the TeV into a dry squeeze at 12:20 PM.

At 12:44 PM, Operations put the TeV into shot setup and at 3:14 PM established store 3191. Its initial luminosity was 39.8E30.

Another Booster power supply (EAPS) tripped off at 3:19 PM with no indication of a problem. Operators contacted experts who came in to make repairs.

Operators made Pbar ready for an access at 3:33 PM. Experts reported finding a bad connector on D:ISEPV. They began repairs at 4 PM.

The Booster experts completed their repairs and Operations established beam to MiniBooNE at 4:46 PM.

At 5:55 PM, the Pbar experts finished their repairs and stacking resumed.

Monday January 26
The midnight shift began with Operations monitoring store 3191, with Pbar stacking, and with MiniBooNE taking beam.

MI suffered a vacuum pump trip in its RF sector. It reset okay.

Today's Plans
The plan for today is to stack and store. The Collider Detector Facility will be allowed to access their collision hall between stores.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/26/2004   email Fermilab